OpenStudy (anonymous):

Please help? I need an answer to this question, because all the ones I have gotten before are wrong. :(

OpenStudy (anonymous):

where is the question?

OpenStudy (anonymous):

i put the wrong attachment on hang on

OpenStudy (id21):

\[\sqrt[4]{81a^8b^{10}}=\sqrt[4]{3^4a^8b^{10}}=3^{\frac{4}{4}}a^{\frac{8}{4}}b^{\frac{10}{4}}=3a^2b^{\frac{5}{2}}=3a^2b^{\frac{4}{2}+\frac{1}{2}}=3a^2b^2\sqrt b\]

OpenStudy (anonymous):

Oh my gosh! Thank you so much.